# Learning How to Vote With Principles Axiomatic Insights Into the Collective Decisions of Neural Networks

### v1

@@ 6 Results and Analysis
6.1 Correct for the Right Reasons?
Recent work in computer science has studied the capabilities of neural networks to learn voting rules (Anil and Bao,, 2021; Burka et al.,, 2022), but without asking whether “the system performs well for the right reasons” (Bender and Koller,, 2020, p. 5192). Here we use voting-theoretic axioms to shed light on the learning behavior of neural networks, specifically aiming to distinguish solely accurate versus principled learning.
@@ 6.1 Correct for the Right Reasons?
We train each one of the three neural network architectures (MLP, CNN, and WEC) on data from each one of the three basic voting rules (Plurality, Borda, and Copeland) using four different sampling distributions (IC, Urn, Mallows, and Euclidean). We report the results as relative accuracy and axiom satisfaction, i.e.,
@@
For example, if the model has $95\%$ identity accuracy, then, since the rule trivially has $100\%$ accuracy, the relative identity accuracy is $100\%-95\%=5\%$ (i.e., the error). If the model has $35\%$ satisfaction of the independence axiom and the rule only $30\%$, then the relative independence satisfaction is $30\%-35\%=-5\%$.
@@ Results.
Figure 4: Training the three architectures (MLP, CNN, and WEC) on data from Plurality, Borda, and Copeland (the three bars in each plot) with IC samples and comparing the errors in both accuracy and axiom satisfaction.
Notably, across all voting rules, architectures, and distributions, we see large violations of neutrality despite high accuracy (e.g., $4.6\%$ identity-accuracy error for the WEC architecture when trained on the Plurality rule but still $19.5\%$ neutrality error). Large violations of anonymity are also observed under the MLP and CNN architectures (the WEC is anonymous by design). This is particularly noteworthy since anonymity and neutrality are always satisfied by the given voting rules. The MLP and CNN models regularly violate anonymity more than neutrality (with the models trained on Plurality demonstrating the smallest such difference).
Regarding the other axioms, all models adhere perfectly to Pareto, in accordance with the voting rules on which they are trained. The MLP and WEC models trained on Plurality seem to satisfy the Condorcet principle more than Plurality does, but the opposite holds for the CNN model. Along a similar line, the MLP model trained on Borda seems to satisfy the Condorcet principle more than Borda does, but this is not the case for the CNN and WEC models. Since Copeland always satisfies the Condorcet principle, the corresponding principled error of all models is positive—yet, it is rather small. The MLP and WEC models trained on Plurality and Borda, as well as the CNN model trained on Plurality, satisfy independence to a similar degree as the rules do on which they are trained. All models trained on Copeland satisfy independence more than Copeland does, and the same holds for the CNN model trained on Borda.
@@ Discussion.
Regarding the learnability of different voting rules, the simplicity of Plurality is probably the reason behind the high accuracy with which all models learn it. However, this simplicity also renders Plurality problematic in other contexts (Laslier,, 2011).
The models take a stance on the well-documented tension between anonymity and neutrality. [^7] No voting rule that always elects a single winner can simultaneously uphold both anonymity and neutrality. They tend to favor outcomes that align more closely with the former than with the latter. This inclination exposes an inherent bias within neural networks when navigating fundamental democratic axioms.
As the architectures are not invariant to permutations of the input data, the severe violations neutrality (and of anonymity for MLPs and CNNs) are not a priori surprising. What is surprising is that this violation persists even for high accuracy with respect to rules that are perfectly neutral and anonymous.
Overall, our experiment on accurate versus principled learning within voting contexts highlights precisely the importance of the reasons behind automated decision-making. Outcomes that are mimicking well-defined voting rules are arguably still unsafe to rely on, since they do not come with a guarantee of respecting the principles on which those rules are built.
6.2 Learning Principles by Example?
Can we teach neural networks voting-theoretic principles, beyond merely presenting data from various voting rules? A natural approach for integrating expert knowledge in neural networks is data augmentation. In the voting-context, this was proposed by Xia, (2013) but has not been tested in practice, to the best of our knowledge. We focus on the neutrality axiom, since it was violated most, and we also test the effects of data augmentation on the model’s accuracy.
@@ 6.2 Learning Principles by Example?
We train each one of the three neural network architectures (MLP, CNN, and WEC) on data from each one of the three basic voting rules (Plurality, Borda, and Copeland); but we vary the ratio $p$ of sampled data and augmented data, while keeping the total number of data points fixed. Thereby, any improvement comes from the quality of the augmented data points and not from merely a higher quantity of data points.
Specifically, given a percentage $1\leq p\leq 100$, we first sample $p$ (times the total data size) many profiles and compute the corresponding winning sets according to the considered rule; call these the sampled data points. Then we generate the remaining $100-p$ many data points as follows: we randomly pick one of the sampled data points $(\boldsymbol{P},S)$ and a permutation $\sigma$ of the alternatives, and then add the data point $(\sigma(\boldsymbol{P}),\sigma(S))$; call these the augmented data points. Thus, the augmented data points are “neutrality variations” of the sampled data points. For different choices of $p$, we then test the models’ neutrality satisfaction and accuracy.
Results.
Results for IC sampling are exhibited in Figure 5 (and for the Mallows model in Figure 8 in the Appendix). We find that data augmentation does not improve adherence to neutrality: the ratio $p$ between sampled and augmented data does not seem to correlate with neutrality satisfaction. For $p<10\%$, i.e., with almost only augmented data, both accuracy and neutrality satisfaction are unsatisfactory, so data augmentation only becomes relevant for $p\geq 10\%$. Here accuracy is stable: it does not vary by more than $5\%$. In some cases, neutrality is equally stable: for the CNN on all rules and the MLP on Borda (certainly for $p\geq 25\%$, with slightly worse neutrality satisfaction for smaller $p$). In the remaining non-stable cases, the best neutrality satisfaction is achieved for $p=100\%$, i.e., without augmented data—with only negligible exceptions. [^8] The only two exceptions are the CNN on Plurality (where neutrality is most satisfied at $p=75\%$ but to a very similar degree as for $p=100\%$) and the CNN on Copeland (where neutrality is minimized at $p=25\%$). Moreover, CNN on Borda and MLP on Copeland have a local—albeit not global—minimum at $p=25\%$. Thus, while there might be some special cases where neutrality is improved in the highly augmented scenario, this is not enough to consider data augmentation as a successful strategy to improve neutrality satisfaction (which is what we are concerned with here). Thus, neither in the stable nor the unstable cases can we see reliable comparative improvements in neutrality satisfaction with more neutrality-augmented data.
[…]
Figure 5: For each architecture, the accuracy and neutrality error across different ratios of augmented data. For example, $10\%$ in the $x$ axis means that the dataset consists of $10\%$ sampled data and $90\%$ augmented data.
Discussion.
Learning voting-theoretic principles by examples—augmented to the training data—does not seem to work for neural networks: Comparatively more neutrality-augmented data does not lead to higher neutrality satisfaction. However, an advantage of data augmentation is a drastic increase in data efficiency when we only aim for accuracy. Sampling only $10\%$ of the total data set (and using neutrality augmented data for the remaining $90\%$) does not substantially decrease the MLP’s or WEC’s accuracy in comparison to sampling the whole data set. This is crucial if we use real and not sampled election data, where having access to a vast amount of data points is practically impossible. Even when more data is needed to increase the accuracy of network, we could build an appropriate data set based on a limited amount of real data points and then augment it via the neutrality axiom.
6.3 Rule Synthesis Guided by Principles?
We saw that neural networks, when trained on data from established voting rules, struggle to vote with principles. This raises the question: can we directly train neural networks to form principled collective decisions, without relying on any pre-existing voting rules? This will be limited by Arrow’s Impossibility Theorem (Arrow,, 1951): a voting rule cannot simultaneously satisfy anonymity, Pareto, and independence. Neural network-based approaches also face this impossibility. However, how close can we get to full axiom satisfaction? We design an optimization task, using custom loss functions, to guide neural networks in learning novel and principled voting rules.
Design.
@@ 6.3 Rule Synthesis Guided by Principles?
We train each one of the three neural network architectures (MLP, CNN, and WEC) on the loss functions defined in Section 4.4 which represent the axioms anonymity, neutrality, Condorcet, Pareto, and independence. Since neural networks could attempt to vacuously satisfy the axioms by proposing no winner, we also consider the “No-winner” loss function, which demands the winning sets to be nonempty. Moreover, by Arrow’s Theorem, the axioms cannot be jointly satisfied and will, hence, negatively influence each other. So optimizing for all axioms is not necessarily the best. Instead, we pick, for each architecture, a set $\mathcal{O}$ of objectives that we optimize for. For WEC, we choose: no winner, Condorcet, and Pareto. For MLP and CNN, we add: anonymity and independence. Then the optimization problem is:
@@
where the loss functions $L_{O}$ are described in Section 4.4 and $D$ is the chosen distribution of profiles $\mathbf{P}$ (IC or Mallows). Note that, unlike the previous experiments, this is an unsupervised learning task.
In order to have an architecture that also is neutral by design (not just anonymous by design like the WEC), we design a further decoding function in addition to the one used so far (Section 4.3). This neutrality-averaged decoding works as follows (cf. Burka et al.,, 2022). Given an input profile, we first generate all alternative-permuted versions of the profile, then compute the logits-predictions of the model on each of those permuted profiles (in one batch), next de-permute the predictions again and average all of them, and finally we turn those average logits into a winning set with the decoding function used so far.
@@ Results.
Table 1 shows the axiom satisfaction of different neural networks (bottom) and, for comparison, of several known rules from voting theory (top), all using IC sampling (for Mallows see Table 6 in the Appendix). The best neural-network based rule in terms of axiom satisfaction is the neutrality-averaged WEC, with close contestants the neutrality-averaged CNN and MLP. The neutrality-averaged WEC clearly outperforms the classic Plurality, Borda, and Copeland rules in every single axiom. Even when we consider more modern rules in voting theory, the neutrality-averaged WEC is competitive: the existing rule with highest axiom satisfaction is Stable Voting and its edge is marginal, with its average axiom satisfaction being less than $1\%$ higher than that of the neutrality-averaged WEC. [^9] Table 3 in the Appendix suggests that more gradient steps do not further improve the results. (In the Mallows case, the neutrality-averaged WEC even just beats all other voting rules, see table 6 in the Appendix. [^10] Tables 4 and 5 in the Appendix suggest the statistical robustness of these results by reporting the average results across 5 runs of this experiment with different seeds.)
@@
Table 1: Axiom satisfaction of different rules (top part of the table) and models (bottom part of the table), for IC sampling. Rounded to one decimal. The names of the models are explained as follows: The letters after the architecture type indicate how the voting rule is computed from the model: p–plain (i.e., no averaging), n–neutrality-averaged, na–neutrality-and-anonymity-averaged. The letters in the brackets indicate which axioms the model optimized for during training: NW–No winner, A–Anonymity, C–Condorcet, P–Pareto, I–Independence. All models have been trained for 15k gradient steps with batch size 200 and the same seed. The WEC by far was the fastest.
In addition to examining axiom satisfaction, we should also consider how often the examined rules produce the same outcomes: because similar axiom satisfaction does not imply similarity of outcomes. [^11] For example, the Blacks and Weak Nanson rules are close in average axiom satisfaction (less than $1\%$ difference), but Table 2 shows that more than $8\%$ of the time they propose a different set of winners. Table 2 describes similarity in outcome between the five rules which excelled in axiom satisfaction, using IC sampling (for Mallows, see Table 7 in the Appendix). In particular, we see that the rule discovered by the neutrality-averaged WEC model is substantially different from the existing voting rules: it proposes different outcomes than each one of them at least $9.3\%$ of the time. In comparison, Stable Voting that was found best in Table 1 disagrees with Borda and Copeland $8.9\%$ of the time and with Weak Nanson and Blacks only $6.6\%$ of the time. Thus, the discovered rule not only is competitive in axiom satisfaction, it also is novel, i.e., substantially different from existing voting rules.
WEC n
Blacks
Stable Voting
Borda
Weak Nanson
Copeland
WEC n 100 90.7 90.2 89.6 88.1 87.4
Blacks 100 95.65 95.45 91.82 90.56
Stable Voting 100 91.1 93.38 91.99
Borda 100 87.27 86.01
Weak Nanson 100 92.26
@@
Table 2: Similarities between the rules. Computed on 10.000 IC-sampled profiles. For example, the entry $90.2$ in row ‘WEC n’ and column ‘Stable Voting’ means that in $90.2\%$ of the sampled profiles Stable Voting outputs the same winning set as the neutrality-averaged WEC.
To illustrate the difference between the discovered and the existing rules, Figure 6 shows an example of a profile where the winning set provided by the neutrality-averaged WEC model is different to all the winning sets provided by the considered existing voting rules. The choice of the WEC also has intuitive plausibility: it chooses alternative $a$ which, among the eight voters, is three times the most preferred option and two times the second-most preferred option.
@@
$\{a\}$ neutrality-averaged WEC
@@
Figure 6: Profile where the ‘WEC n’ model disagrees with existing voting rules. The winning sets for each rule are mentioned below the table.

### v2

@@ 6. Results and Analysis
6.1. Experiment 1: Correct for the Right Reasons?
Recent work in computer science has studied the capabilities of neural networks to learn voting rules (Anil and Bao, 2021; Burka et al., 2022), but without asking whether “the system performs well for the right reasons” (Bender and Koller, 2020, p. 5192). Here we use voting-theoretic axioms to shed light on the learning behavior of neural networks, specifically aiming to distinguish solely accurate versus principled learning.
@@ 6.1. Experiment 1: Correct for the Right Reasons?
We train each of the three neural network architectures (MLP, CNN, and WEC) on data from each one of the three basic voting rules (Plurality, Borda, and Copeland) using four different sampling distributions (IC, Urn, Mallows, and Euclidean) with the parameters mentioned in Section 5.1. We report the results as relative accuracy and axiom satisfaction, i.e.,
@@
For example, if the model has $95\%$ accuracy, then, since the rule has $100\%$ accuracy, there is a relative accuracy loss of $100\%-95\%=5\%$. If the model has $35\%$ satisfaction of the independence axiom and the rule only $30\%$, then the relative independence satisfaction is $30\%-35\%=-5\%$, so there is a relative independence gain of $5\%$.
@@ Results.
Three barplots describing the accuracies and axiom satisfactions of the three architectures.
Figure 4. Training the three architectures (MLP, CNN, and WEC) on data from Plurality, Borda, and Copeland (the three bars in each plot) with IC samples and comparing the errors in both accuracy and axiom satisfaction. The error describes the rule’s evaluation minus the model’s evaluation. Intuitively, $2.8\%$ accuracy error means $2.8\%$ loss in accuracy: the rule by definition is correct so it has $100\%$ accuracy, but the model obtains only $97.2\%$ accuracy; similarly, $6.5\%$ Condorcet error means $6.5\%$ gain in accuracy: the rule has $73.25\%$ Condorcet satisfaction, but the model obtains $79.75\%$ Condorcet satisfaction.
Notably, across all voting rules, architectures, and distributions, we see large losses in neutrality despite only low losses in accuracy (e.g., $4.6\%$ relative identity-accuracy loss but $19.5\%$ relative neutrality loss for the WEC architecture when trained on the Plurality rule). Large anonymity losses are also observed under the MLP and CNN architectures (the WEC is anonymous by design). This is particularly noteworthy since anonymity and neutrality are $100\%$ satisfied by the given voting rules. The MLP and CNN models regularly show larger neutrality losses than anonymity losses (with the models trained on Plurality demonstrating the smallest such difference).
Regarding the other axioms, all models adhere perfectly to Pareto, in accordance with the voting rules on which they are trained. The MLP and WEC models trained on Plurality seem to exhibit relative Condorcet gains, but Condorcet losses are found for the CNN model. Along a similar line, the MLP model trained on Borda obtains relative Condorcet gains, but this is not the case for the CNN and WEC models. Since Copeland always satisfies the Condorcet principle, there is a relative Condorcet loss for all models—yet, it is rather small. The MLP and WEC models trained on Plurality and Borda, as well as the CNN model trained on Plurality, satisfy independence to a similar degree as the rules do on which they are trained. All models trained on Copeland exhibit relative independence gains, and the same holds for the CNN model trained on Borda.
Discussion.
Regarding the learnability of different voting rules, the simplicity of Plurality is probably the reason behind the high accuracy with which all models learn it. However, this simplicity also renders Plurality problematic in other contexts (Laslier, 2011).
The models take a stance on the well-documented tension between anonymity and neutrality. [^16] No voting rule that always elects a single winner can simultaneously be anonymous and neutral. They tend to favor outcomes that align more closely with the former than with the latter. This inclination exposes an inherent bias within neural networks when navigating fundamental democratic axioms.
As the architectures are not invariant to permutations of the input data, the severe violations of neutrality (and of anonymity for MLPs and CNNs) are not a priori surprising. However, these violations persist even for high accuracy with respect to rules that are perfectly neutral and anonymous.171717 Actually, whether this is surprising depends on which of the following two intuitions one has. The first intuition regards these results as surprisingly bad: Given the high accuracy, we may expect that the neural network should have ‘gotten the idea’ of the voting rule, and hence of its anonymity and neutrality, so the amount of violations is surprising. The second intuition regards the results as surprisingly good: For anonymity and, respectively, neutrality to be satisfied on a given profile, we require the neural network to output the correct answer on 50 permutations of the profile, while accuracy requires being correct only on that very profile, so it is not surprising that the neural network struggles more with anonymity and neutrality. But whether surprising or not, the results stay the same: for the specified percentages of considered profiles, we have violations in anonymity and neutrality, i.e., we do not have high certainty (at least one failure in 50 checks) that the neural network outputs the desired answer under permutations.
Overall, our experiment on accurate versus principled learning highlights the importance of the reasons behind automated decision-making. Outcomes that mimic well-defined voting rules are arguably still unreliable, since they do not come with a guarantee of respecting the principles on which those rules are built.
6.2. Experiment 2: Learning Principles by Example?
Can we teach neural networks voting-theoretic principles, beyond merely presenting data from various voting rules? A natural approach to integrate expert knowledge in neural networks is data augmentation. In the voting context, this was proposed by Xia (2013) but has not been tested in practice, to the best of our knowledge. We focus on the anonymity and neutrality axioms, since they were violated most, and we also test the effects of data augmentation on the model’s accuracy.
Asking if data augmentation helps can be understood in two ways: First, does training with augmented data increase axiom satisfaction without diminishing accuracy? Second, if so, does it do this better than just training with sampled data points? A ‘yes’ to the first question improves data efficiency: we get at least as good a model even when only part of the data is ‘real’ and the rest is augmented. A ‘yes’ to the second question means that we can actually improve our model’s performance in the preceding experiment.
Design.
We test data augmentation in two versions. In the first version, we form an initial dataset (i.e., pairs of a sampled profile with corresponding winning set) and we train an architecture on this dataset. Then we continue training it on augmented data points, i.e., data points obtained from the initial data points by renaming alternatives (‘neutrality variations’) or by renaming voters (‘anonymity variations’). For comparison, we make a copy of our model after the initial training and continue training the copy with the same number, but sampled data points (rather than augmented data points). If the model trained on augmented data improves axiom satisfaction without worsening accuracy, we get a ‘yes’ to the first question. If it also is better than the copied model, we also get a ‘yes’ to the second question. Otherwise, any improvement only comes from a mere increase in the quantity of data points and not from their quality.
A potential issue of this version is that, during the continued training with augmented data, the model does not see any further sampled data and hence might lose in accuracy. The second version fixes this issue by making sure that each training batch consists of $p$ percent sampled data points with the remaining data points being neutrality variations (resp., anonymity variations) of those sampled data points. For different choices of $p$, we then test the models’ achieved axiom satisfaction and accuracy. We get a ‘yes’ to the first (resp., second) question if axiom satisfaction and accuracy are not worse (resp., better) for lower values of $p$ when compared to $p=100\%$ (i.e., only sampled data).
Results.
Results for IC-sampling and neutrality augmentation are exhibited in Figure 5. (Appendix C presents results also for other distributions and anonymity augmentation.) Overall, we find a ‘yes’ to the first question but a ‘no’ to the second: training with augmented data does not hurt accuracy, but it does not reliably improve axiom satisfaction. In the first version of the experiment depicted at the top row and the left plot of the second row in Figure 5, augmented data does not seem to be advantageous for neutrality relatively to sampled data (on the contrary, it often seems harmful, e.g., when applying CNN or WEC); in the second version of the experiment depicted at the bottom row and the right plot of the second row in Figure 5, the ratio $p$ between sampled and augmented data does not seem to correlate with neutrality satisfaction either.
@@
Regarding the second version, when $p<10\%$, i.e., with almost only augmented data, both accuracy and neutrality satisfaction are unsatisfactory, so data augmentation only becomes relevant for $p\geq 10\%$. Here accuracy is stable: it does not vary by more than $5\%$. In some cases, neutrality is equally stable: for the CNN on all rules and the MLP on Borda (certainly for $p\geq 25\%$, with slightly worse neutrality satisfaction for smaller $p$). In the remaining non-stable cases, the best neutrality satisfaction is achieved for $p=100\%$, i.e., without augmented data—with only negligible exceptions. [^18] The only two exceptions are the CNN on Plurality (where neutrality is most satisfied at $p=75\%$ but to a very similar degree as for $p=100\%$) and the CNN on Copeland (where neutrality is minimized at $p=25\%$). Moreover, CNN on Borda and MLP on Copeland have a local—albeit not global—minimum at $p=25\%$. Thus, while there might be some special cases where neutrality is improved in the highly augmented scenario, this is not enough to consider data augmentation as a successful strategy to improve neutrality satisfaction (which is what we are concerned with here). Thus, neither in the stable nor the unstable cases can we see reliable comparative improvements in neutrality satisfaction with more neutrality augmented data.
@@
Six plots describing the results of the second experiment.
Figure 5. Top row and second row on the left: The first version. Pretraining on IC-sampled data from Plurality, continuing with neutrality variations and comparing this with identity accuracy to sampled data. Rest: The second version.
@@ Discussion.
Learning voting-theoretic principles by examples—augmented to the training data—does not seem to work for neural networks: Comparatively more neutrality-augmented or anonymity-augmented data does not necessarily lead to higher neutrality or anonymity satisfaction. However, an advantage of data augmentation is a drastic increase in data efficiency when we only aim for accuracy. For instance, sampling only $10\%$ of the total data set (and using neutrality augmented data for the remaining $90\%$) does not substantially decrease the MLP’s or WEC’s accuracy in comparison to sampling the whole data set. This is crucial if we use real and not sampled election data, where having access to a vast amount of data points is practically impossible. Even when more data is needed to increase the accuracy of network, we could build an appropriate data set based on a limited amount of real data points and then augment it via the neutrality axiom. This gives us an answer to the two questions raised earlier on: ‘Yes’, training with additional, augmented data points can increase axiom satisfaction, but ‘no’, not better than just training with equally many sampled data points.
6.3. Experiment 3: Rule Synthesis Guided by Principles?
We saw that neural networks, when trained on data from established voting rules, struggle to vote with principles. This raises the question: can we directly train neural networks to form principled collective decisions, without relying on any pre-existing voting rules? This will be limited by Arrow’s Impossibility Theorem (Arrow, 1951): a voting rule cannot simultaneously satisfy anonymity, Pareto, and independence. Neural network-based approaches also face this impossibility. However, how close can we get to full axiom satisfaction? We design an optimization task, using custom loss functions, to guide neural networks in learning novel and principled voting rules.
@@ 6.3. Experiment 3: Rule Synthesis Guided by Principles?
We train each one of the three neural network architectures (MLP, CNN, and WEC) on the loss functions defined in Section 4.4 which represent the axioms anonymity, neutrality, Condorcet, Pareto, and independence. Since neural networks could attempt to vacuously satisfy the axioms by proposing no winner, we also consider the “No-winner” loss function, which demands the winning sets to be nonempty. Moreover, by Arrow’s Theorem, the axioms cannot be jointly satisfied and will, hence, negatively influence each other. So optimizing for all axioms is not necessarily the best. We pick, for each architecture, a set $\mathcal{O}$ of objectives that we optimize for. For WEC, we choose: no winner, Condorcet, and Pareto. (Appendix D.5 establishes in an ablation study the optimality of this choice.) For MLP and CNN, we add: anonymity and independence. Then the optimization problem is:
@@
where the loss functions $L_{O}$ are described in Section 4.4 and $D$ is the chosen distribution of profiles $\mathbf{P}$. Note that, unlike the previous experiments, this is an unsupervised learning task.
In order to have an architecture that is also neutral by design (not just anonymous by design like the WEC), we design a further decoding function in addition to the one used so far (Section 4.3). This neutrality-averaged decoding works as follows (cf. Burka et al., 2022). Given an input profile, we first generate all alternative-permuted versions of the profile, then compute the logits-predictions of the model on each of those permuted profiles (in one batch), next de-permute the predictions again and average all of them, and finally we turn those average logits into a winning set with the decoding function used so far.
@@ Results.
Table 1 shows the axiom satisfaction of different neural networks (bottom) and, for comparison, of several known rules from voting theory (top), all using IC sampling. (Appendix D shows the results for the other distributions. See also Figure 16 in the Appendix for the evolution of the loss during optimization.)
The best neural-network based rule in terms of axiom satisfaction is always the neutrality-averaged WEC, with close contestants the neutrality-averaged CNN and MLP. The neutrality-averaged WEC outperforms the classic Plurality, Borda, and Copeland rules in every single axiom, except for a slight loss on Condorcet. Even when we consider more modern rules in voting theory, the neutrality-averaged WEC is competitive: the existing rule with highest axiom satisfaction is Stable Voting and its edge is marginal, with its average axiom satisfaction being less than $1\%$ higher than that of the neutrality-averaged WEC. [^19] Table 4 in the Appendix suggests that more gradient steps do not further improve the results. In fact, when averaging five runs of checking axiom satisfaction (which always involves some stochasticity), the neutrality-averaged WEC even comes out better than the rules: see Table 5 in the Appendix. (This is also true for Euclidean but not for Mallows and Urn.) In any case, the neutrality-averaged WEC has a comparably good axiom satisfaction as the best voting rules known today.
@@
Table 1. Axiom satisfaction of different rules (top part of the table) and models (bottom part of the table), for IC sampling. Rounded to one decimal. The names of the models are explained as follows: The letters after the architecture type indicate how the voting rule is computed from the model: p–plain (i.e., no averaging), n–neutrality-averaged, na–neutrality-and-anonymity-averaged. The letters in the brackets indicate which axioms the model optimized for during training: NW–No winner, A–Anonymity, C–Condorcet, P–Pareto, I–Independence. All models have been trained for 15k gradient steps with batch size 200.
In addition to examining axiom satisfaction, we should also consider how often the examined rules produce the same outcomes: because similar axiom satisfaction does not imply similarity of outcomes. [^20] For example, the Blacks and Weak Nanson rules are close in average axiom satisfaction (less than $1\%$ difference), but Table 2 shows that more than $8\%$ of the time they propose a different set of winners. Table 2 describes similarity in outcome compared to the five closest rules, using IC sampling (again, see Appendix D for the other distributions). In particular, we see that the rule discovered by the neutrality-averaged WEC model is substantially different from the existing voting rules: it proposes different outcomes than each one of them, according to identity accuracy, at least $9.3\%$ of the time (resp., $10.6\%$ for Mallows, $11.1\%$ for Urn, and $7.8\%$ for Euclidean, see Tables 7, 10, and 13 in the Appendix). In comparison, Stable Voting, which was found best in Table 1, disagrees with Borda and Copeland $8.9\%$ of the time and with Weak Nanson and Blacks only $6.6\%$ of the time. Thus, the discovered rule not only is competitive in axiom satisfaction, it also is novel, i.e., substantially different from existing voting rules.
Moving from hard to soft accuracy, the neutrality-averaged WEC produces winning sets that are a subset (resp., superset) of those of existing rules at least $95.4\%$ (resp., $95.3\%$) of the time for IC, and similarly for other distributions. This means that even if the results of the model differ from those of existing voting rules, very frequently they do so by only excluding or by only adding certain winning alternatives. This is not unique to our ML model—it is also the case between known voting rules that exhibit high axiomatic satisfaction: for example, the outcome of Stable Voting is a subset (resp., superset) of the outcome of Weak Nanson $97.9\%$ (resp., $94.35\%$) of the time.
Identity accuracy WEC n Blacks Stable Voting Borda Weak Nanson Copeland
WEC n 100 91 90.5 89.5 88.4 87.7
Blacks 100 95.71 95.13 91.26 90.57
Stable Voting 100 90.84 93.5 91.67
Borda 100 86.39 85.7
Weak Nanson 100 92.43
@@
Subset accuracy WEC n Blacks Stable Voting Borda Weak Nanson Copeland
WEC n 100 93.6 92.7 93.9 93 95.3
Blacks 95.6 100 96.53 97.3 95.57 98.2
Stable Voting 95.4 97.21 100 94.51 97.9 99.59
Borda 93.8 95.13 91.66 100 90.7 93.33
Weak Nanson 92 92.69 94.35 89.99 100 97.71
Copeland 90.4 91.29 91.73 88.59 94.15 100
Table 2. Similarities between the rules. Computed on 10,000 IC-sampled profiles. For example, in the top table (‘identity accuracy’), the entry $90.5$ in row ‘WEC n’ and column ‘Stable Voting’ means that, in $90.5\%$ of the sampled profiles, Stable Voting outputs the identical winning set as the neutrality-averaged WEC. Hence the values below the diagonal are symmetric and thus omitted. In the bottom table (‘subset accuracy’), the entries show when the rule in the row outputs a winning set that is a subset of the rule in the column. So the entry $92.7$ in row ‘WEC n’ and column ‘Stable Voting’ means that, in $92.7\%$ of the sampled profiles, the winning set outputted by the neutrality-averaged WEC is a subset of the winning set outputted by Stable Voting. This table is not symmetric, because the entry $95.4$ in row ‘Stable Voting’ and column ‘WEC n’ means that, in $95.4\%$ of the sampled profiles, the winning set outputted by Stable Voting is a subset of the model’s winning set (equivalently, the model’s winning set is a superset of the rule’s winning set).
To illustrate the difference between the discovered and the existing rules, Figure 6 shows an example of a profile where the winning set provided by the neutrality-averaged WEC is different to all the winning sets provided by the considered existing voting rules. The choice of the WEC also has intuitive plausibility: it chooses alternative $a$ which, among the eight voters, is three times the most preferred option and two times the second-most preferred option. The sigmoids indicate that alternative $b$ was also a close competitor for the winning set, and would indeed win under many of the known rules from voting theory.
@@
$\{a\}$ neutrality-averaged WEC, with sigmoids (rounded) aa:.51, bb:.49, cc:.31, dd:.32, ee:.43
@@
[…]
A profile on which the ‘WEC n’ model disagress with existing voting rules.
Figure 6. Profile where the ‘WEC n’ model disagrees with existing voting rules. The winning sets for each rule are mentioned below the table.